OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. 9 * 72 = 9 * ? + 9 * 2 =?

OpenStudy (hhmorris98):

9*72=9*70+9*2=648

OpenStudy (anonymous):

ohooo well that is easier than i had thought thanks for the help bro! @hhmorris98

OpenStudy (hhmorris98):

Np I gotcha whenever you need it.